please help me solve for x and y
it takes bob the boy scout 10 hours to paddle upstream {against the current} a distance 15 miles. when he turns around he finds it only takes him 5 hour to paddle 22.5 miles with the current.how fast is bob's boat going in stilll water and what is the speed of the curent of the river?
please solve this problem right now!!!! thank you

Answer by drk(1908)

This is a RTD problem. Here is a table based on given information
paddle . . . . . . . . .rate . . . . . . . . . time . . . . . . . . . distance
up stream . . . . . .r - c . . . . . . . . . . .10 . . . . . . . . . . 10r - 10c =15
down stream . . . .r + c . . . . . . . . . . .5 . . . . . . . . . . .5r + 5c = 22.5
totals . . . . . . . . .
Looking at column 3 we get
(i) 10r - 10c = 15
(ii) 5r + 5c = 22.5
multiply (ii) by 2 to get
(iii) 10r + 10c = 45
add (i) and (iii) to get
(iv) 15r = 60
divide to get
(v) r = 4
This means that the current is
c = 2.5